# Analyzing User Characteristics of Hate Speech Spreaders on Social Media

### v2

@@ Abstract.
Hate speech on social media threatens the mental and physical well-being of individuals and contributes to real-world violence. Resharing is an important driver behind the spread of hate speech on social media. Yet, little is known about who reshares hate speech and what their characteristics are. In this paper, we analyze the role of user characteristics in hate speech resharing across different types of hate speech (e.g., political hate). For this, we proceed as follows: First, we cluster hate speech posts using large language models to identify different types of hate speech. Then we model the effects of user attributes on users’ probability to reshare hate speech using an explainable machine learning model. To do so, we apply debiasing to control for selection bias in our observational social media data and further control for the latent vulnerability of users to hate speech. We find that, all else equal, users with fewer followers, fewer friends, fewer posts, and older accounts share more hate speech. This shows that users with little social influence tend to share more hate speech. Further, we find substantial heterogeneity across different types of hate speech. For example, racist and misogynistic hate is spread mostly by users with little social influence. In contrast, political anti-Trump and anti-right-wing hate is reshared by users with larger social influence. Overall, understanding the factors that drive users to share hate speech is crucial for detecting individuals at risk of engaging in harmful behavior and for designing effective mitigation strategies.
@@ 1. Introduction
Online hate speech poses a significant problem to society. Around one in four U.S. adults has experienced online hate (Vogels, 2021), and, in particular, young adults suffer from the psychological consequences (Saha et al., 2019). Hate speech not only threatens users’ mental and physical well-being but has also the potential to fuel radicalization and incite real-world violence (Müller and Schwarz, 2021), as was the case for the suspect of the 2018 Pittsburgh synagogue shooting (Roose, 2018).
@@
On social media, reshares play a vital role in disseminating content beyond immediate circles (Subbian et al., 2017). One risk of reshares is that harmful content can go viral, such as propaganda (Geissler et al., 2023), fake news (Cheng et al., 2021; Shao et al., 2018; Pröllochs et al., 2021a, b), and content from untrustworthy sources (Guess et al., 2023). Prior research found that reshares also make hate speech go disproportionally more viral than normal content (Maarouf et al., 2024; Mathew et al., 2019), which is exploited by hateful users to reach more people (Goel et al., 2023).
@@
However, prior work has mostly looked into who posts hate speech (Hua et al., 2020; Ribeiro et al., 2018) and who is targeted by it (ElSherief et al., 2018). Hateful users are characterized by high activity levels, complex word usage, and high connectivity among each other (Hua et al., 2020; Ribeiro et al., 2018). Similarly, targets of hate speech show high levels of activity, but they also tend to be older accounts and have more followers (ElSherief et al., 2018). However, reshares help hate speech reach larger audiences and are an important driver to the virality of hate speech (Maarouf et al., 2024). Nevertheless, no work has been conducted to profile the users who reshare hate speech. Hence, in this work, we analyze which user attributes (such as, e.g., verified status, account age, or number of posts, followers, and friends) make users reshare hate speech while taking into account the different types of hate speech.
@@
Estimating the effects of user attributes on users’ probability to reshare hate speech is difficult. One challenge is the use of observational social media data to estimate the effects. Observational data from social media is likely biased because it is subject to selection bias from the user and the recommendations from online platforms (Schnabel et al., 2016). For example, users only interact with content that they like and online platforms only recommend content that they assume a user is interested in (selection bias). We later employ inverse propensity scoring (IPS) (Rosenbaum and Rubin, 1983) to reweight the observed interactions between users and hate speech posts for debiasing. Another reason is that we have to control for the varying vulnerability of users to hate speech as that can cause bias in the estimates of the effects being studied. Previous research argued that past latent vulnerability (PLV) of a user to specific content influences both the user attributes and the probability to reshare (Benevenuto et al., 2009; Cheng et al., 2021). Users may have a different PLV as some users simply share more hate speech than others. Also, the risk of exposure to hate speech is different for each users, e.g. how many items a user has seen hate speech in their feed. In our setting, we control for PLV to hate speech.
@@
In this work, we analyze the role of user characteristics in hate speech resharing across different types of hate speech. Our framework follows three steps: 1 We first cluster hate speech posts into their type using BERTopic (Grootendorst, 2022) and employ automatic cluster labeling with LLAMA-3 (Meta, 2024). This way, we obtain a topic label for each hate speech post. 2 We then compute the likelihood to reshare types of hate speech per user by looking at the proportion of reshares that were from each type. 3 We then model the effects of user attributes on the probability to reshare hate speech while controlling for the PLV. For this, we debias the observational data by adapting inverse propensity scoring (IPS) from causal inference (Rosenbaum and Rubin, 1983), which then reweights the likelihood of users to be exposed to hate speech (Liang et al., 2016; Schnabel et al., 2016). This way, we obtain debiased propensity scores for the users’ vulnerability to hate speech. We then model the debiased PLV of users to hate speech, which we control for in our framework. For this, we learn latent representations of the debiased PLV in the form of debiased embeddings. Then, we model the (non-linear) effects of the user attributes (verified status, account age, number of posts, followers, and friends) on the user’s probability to reshare hate speech in the future using an explainable boosting machine (EBM) (Nori et al., 2019). [^1] Codes and data for reproducibility are available via an anonymous repository for review and will be available in a public repository upon publication: https://anonymous.4open.science/r/analyzing_user_characteristics_hatespeech_spreaders-021C/README.md
@@ 2.1. Hate speech on social media
Another stream of literature conducted user-level analyses. Targets of hate speech tend to be more popular or high profile users (ElSherief et al., 2018), while users that post hate speech are characterized by higher social media activity and tend to act in more connected networks (Hua et al., 2020; Ribeiro et al., 2018). When looking at hate speech in online political discourse, people of color of the Democratic Party, white Republicans, and women in general tend to be the main targets (Solovev and Pröllochs, 2022).
@@ 2.2. Inverse propensity scoring
Click data from observational social media data is subject to selection bias. Specifically, previous research identifies two sources of selection bias: self-selection by the users and recommendations from online platforms (Schnabel et al., 2016). On the one hand, users self-select the content they interact with on social media. On the other hand, online platforms tend to recommend only content to users, that a user is likely interested in. Hence, we must carefully account for selection bias in our observational click data.
@@ 2.3. Modeling past latent vulnerability
We later transfer the idea of recommendation algorithms to model the PLV of users to reshare hate speech. In particular, we employ recommendation algorithms to learn the debiased PLV of users to hate speech from IPS-reweighted data. By doing so, we represent the PLV as a debiased embedding.
@@ 3.1. Overview
Our objective is to model the effects of user attributes $A$ on the users’ probability to reshare different types of hate speech $Y$ while accounting for the user’s (latent) vulnerability $X$. For this, we utilize BERTopic to cluster posts into different types of hate speech and employ a state-of-the-art debiasing framework to model the resharing.
Input: We use the interactions of users with hate speech posts as an input to our framework in the form of a bipartite graph. Let $\mathcal{U}=\{u_{1},\ldots,u_{i}\}$ denote users and =$\mathcal{H}=\{h_{1},\ldots,h_{j}\}$ denote the hate speech posts. The interaction between user $u\in\mathcal{U}$ and hate speech post $h\in\mathcal{H}$ is then a binary variable denoted as $S_{uh}=\{0,1\}$, where $S_{uh}=1$ if user $u$ shared hate speech post $h$, other $S_{uh}=0$.
@@
Clustering hate speech: In the first step, we employ BERTopic (Grootendorst, 2022) combined with HDBSCAN for clustering the hate speech posts $\mathcal{H}$. Subsequently, we label the cluster using automatic labeling with LLAMA-3 (Meta, 2024). Hence, we generate clusters of types of hate speech $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$.
@@
Likelihood to reshare hate speech: In the next step, we compute the likelihood of users to reshare hate speech posts and differentiate between multiple types of hate speech. We write the individual user’s probability to reshare hate speech in general as outcome variable $Y_{u}\in[0,1]$ and the probability to reshare hate speech type $c$ as $Y_{uc}\in[0,1]$.
@@
Modeling hate speech resharing: In the final step, we model the effects of user attributes $A$ on users’ probability to reshare hate speech $Y$ while controlling for PLV ${X}$. For this, we first apply IPS reweighting to debias our observational data, which reweights the propensity $\pi$ of hate speech posts to mimic a randomized trial, and, thereby, we address the issue of selection bias from above. Then, we model the past latent vulnerability to hate speech $X_{u}$ for each user $u$. Lastly, we use an explainable boosting machine (EBM) (Nori et al., 2019) to predict the probability to reshare hate speech. For each user $u$ with $m$ user attributes denoted as vector ${a}_{u}=(a_{1},\ldots,a_{m})$, ${a}_{u}\in A$ and PLV embedding ${X}_{u}$ we model the probability to reshare hate speech $Y_{u}$ ($Y_{uc}$ for different types of hate speech $c$).
@@ 3.2. Clustering hate speech (Step 1)
Given hate speech posts $\mathcal{H}$, we employ BERTopic (Grootendorst, 2022), which leverages BERT embeddings and topic modeling, combined with HDBSCAN for clustering. First, we remove stop words, URL links, and emojis. Then, we generate embeddings for each hate speech root post using the pre-trained BERT model all-MiniLM-L6-v2 to capture the semantic content. Next, we reduce the embedding dimensionality using UMAP to enhance performance and interpretability. Lastly, we apply HDBSCAN to the reduced embeddings to identify clusters $\mathcal{C}$. HDBSCAN is highly flexible as it does not require pre-specified numbers of clusters and it is suitable for noisy data, which makes it an ideal choice for social media data. In addition, HDBSCAN can deal well with clusters that vary in shape and size (Campello, Ricardo J. G. B. et al., 2013).
Subsequently, we label the cluster using automatic labeling with the LLAMA-3 model Llama-3-70b-chat-hf (Meta, 2024). We generate five names per cluster and manually pick the most fitting one by analyzing representative posts. We hereby obtain concise summaries of the primary themes within each cluster, thus providing clear and contextually relevant labels.
3.3. Likelihood to reshare hate speech (Step 2)
(2) $Y_{uc}=\frac{n^{u}_{c}}{n^{u}_{\mathrm{hate}}+n^{u}_{\mathrm{normal}}},$
3.4. Modeling hate speech resharing (Step 3)
@@ 3.4.1. IPS reweighting
Given user group $\mathcal{U}$, the hate speech posts $\mathcal{H}$, and the user-hate speech interactions $S_{uh}$, we apply IPS reweighting to alleviate selection bias and obtain debiased propensity scores $\pi$. The propensity score describes the probability that a user is exposed to hate speech.
Assumption: As in (Joachims et al., 2017), we assume that users share posts when (a) they have been exposed to the post, denoted as binary variable $E_{uh}\in\{0,1\}$ ($E_{uh}=1$ if user $u$ was exposed to hate speech post $h$, $E_{uh}=0$, otherwise), and (b) when they are interested in the post, denoted as binary variable $I_{uh}\in\{0,1\}$ ($I_{uh}=1$ if the user is interested in the post $I_{uh}=0$, otherwise). This yields
(3) $S_{uh}=E_{uh}\cdot I_{uh}.$
We further assume that resharing of hate speech is missing-not-at-random (MNAR) (Little and Rubin, 2020). Hence, the probability of user $u\in\mathcal{U}$ spreading a hate speech post $h\in\mathcal{H}$ is a product of the probability of exposure and the probability of interestingness of a post (Saito, 2020). Formally, we have
@@
Why the propensity score is a biased estimate of $\theta_{uh}$: Simply computing $\pi$ from our data gives a biased estimate for $\theta_{uh}$. The reason is that, when working with observational click data, we face two major challenges: First, we only record positive interactions between users and hate speech posts (i.e., hate speech posts that are reshared by the user). Negative interactions (i.e., hate speech posts that were not reshared by the users) are not recorded. Therefore, we do not know whether (a) the user was not interested in a given post or whether (b) the user was not exposed to the post. Second, the observed interactions are likely subject to selection bias. On the one hand, users are more likely to reshare hate speech posts that are more prevalent. On the other hand, online news platforms are more likely to recommend popular posts to users.
Because of the above, the propensity scores $\pi$ when computed from observational social media data is biased, i.e.,
@@
    where we assume that the probability that a user $u$ is exposed to a hate speech post $h$ is higher if the post has gone more viral (as measured by the number of reshares in the bipartite graph). The IPS for rare events could be very high. Hence, we apply a smoothing operator $\mu=0.5$ for more stable propensity scores (see Appendix C for a robustness check with varying $\mu$).
@@
As a result, we obtain debiased propensity scores. We later perform numerical experiments where we compare the three variants (1)–(3) of our recommendation algorithm to the biased algorithm and find that the debiased recommendation algorithms perform better.
@@ 3.4.2. Modeling past latent vulnerability
Recommender algorithm: In our framework, we employ a Bayesian personalized ranking matrix factorization (BPRMF) (Rendle et al., 2009). We optimize the parameters in the BPRMF using a pairwise Bayesian personalized ranking (BPR) loss (Rendle et al., 2009). By using BPRMF, we address the challenges of missing negative interactions, since it assumes that the observed interactions better explain users’ preferences than unobserved ones and should hence be assigned a higher predictive score. Since we employ three different ways to debias propensity scores, we also obtain three different variants of our debiased recommendation model: BPRMF-V (using $\pi_{\mathrm{virality}}$); BPRMF-F (using $\pi_{\mathrm{follower}}$); and BPRMF-NN (using $\pi_{\mathrm{neural}}$). For each user $u$, we model PLV to hate speech as a 64-dimensional embedding ${X}_{u}$. For comparison, the base model BPRMF is biased since it uses $\pi_{\mathrm{biased}}$ as the propensity score.

### v3

@@ Abstract.
Hate speech on social media threatens the mental and physical well-being of individuals and contributes to real-world violence. Resharing is an important driver behind the spread of hate speech on social media. Yet, little is known about who reshares hate speech and what their characteristics are. In this paper, we analyze the role of user characteristics in hate speech resharing across different types of hate speech (e.g., political hate). For this, we first cluster hate speech posts using large language models into different types of hate speech. Then we model the effects of user attributes on users’ probability to reshare hate speech using an explainable machine learning model. To do so, we apply debiasing to control for selection bias in our observational social media data and further control for the latent vulnerability of users to hate speech. We find that, all else equal, users with fewer followers, fewer friends, fewer posts, and older accounts share more hate speech. This shows that users with little social influence tend to share more hate speech. Further, we find substantial heterogeneity across different types of hate speech. For example, racist and misogynistic hate is spread mostly by users with little social influence. In contrast, political anti-Trump and anti-right-wing hate is reshared by users with larger social influence. Overall, understanding the factors that drive users to share hate speech is crucial for detecting individuals at risk of engaging in harmful behavior and for designing effective mitigation strategies. Disclaimer: This work contains terms that are offensive and hateful. to the nature of the work.
@@ 1. Introduction
Online hate speech poses a significant problem to society. Around one in four U.S. adults has experienced online hate (Vogels, 2021), and, in particular, young adults suffer from the psychological consequences (Saha et al., 2019). Hate speech not only threatens users’ mental and physical well-being but also has the potential to fuel radicalization and incite real-world violence (Müller and Schwarz, 2021), as was the case for the suspect of the 2018 Pittsburgh synagogue shooting (Roose, 2018).
@@
On social media, reshares are vital in disseminating content beyond immediate circles (Subbian et al., 2017). One risk of reshares is that harmful content can go viral, such as propaganda (Geissler et al., 2023), fake news (Cheng et al., 2021; Shao et al., 2018; Pröllochs et al., 2021a, b), and content from untrustworthy sources (Guess et al., 2023). Prior research found that reshares also make hate speech go disproportionally more viral than normal content (Maarouf et al., 2024; Mathew et al., 2019), which is exploited by hateful users to reach more people (Goel et al., 2023).
@@
However, prior work has mostly looked into who posts hate speech (Hua et al., 2020; Ribeiro et al., 2018) and who is targeted by it (ElSherief et al., 2018). Hateful users are characterized by high activity levels, complex word usage, and high connectivity among each other (Hua et al., 2020; Ribeiro et al., 2018). Similarly, targets of hate speech show high levels of activity, but they also tend to be older accounts and have more followers (ElSherief et al., 2018). However, no work has been conducted to profile the users who reshare hate speech, although reshares help hate speech reach larger audiences and are an important driver to the virality of hate speech (Maarouf et al., 2024). Hence, in this work, we analyze the effects of attributes (such as, e.g., verified status, account age, or number of posts, followers, and friends) of users who reshare hate speech while taking into account the different types of hate speech.
@@
Estimating the effects of user attributes on users’ probability to reshare hate speech is difficult. One challenge is the inherent selection bias in observational social media click data, which comes from the users’ self-selection of which content to interact with as well as the recommendations from online platforms (Schnabel et al., 2016). For example, users only interact with content that they choose and online platforms only recommend content that they assume a user is likely to engage with. We later employ inverse propensity scoring (IPS) (Rosenbaum and Rubin, 1983) to reweight the observed interactions between users and hate speech posts for debiasing. Another reason is that we have to control for the varying vulnerability of users to hate speech as that can cause bias in the estimates of the effects being studied. Previous research argued that past latent vulnerability (PLV) of a user to specific content influences both the user attributes and the probability to reshare (Benevenuto et al., 2009; Cheng et al., 2021). Users may have varying PLV as some users simply share more hate speech than others. Also, the risk of exposure to hate speech is different for each user, e.g. how many times a user has seen hate speech in their feed. In our setting, we control for PLV to hate speech.
@@
In this work, we analyze the role of user characteristics in hate speech resharing across different types of hate speech. Our framework follows three steps: 1 We first cluster hate speech posts into their type using BERTopic (Grootendorst, 2022) and employ automatic cluster labeling with LLAMA-3 (Meta, 2024). This way, we obtain a topic label for each hate speech post. 2 We then model the past latent vulnerability of each user to hate speech. For this, we debias the observational data by adapting inverse propensity scoring (IPS) from causal inference (Rosenbaum and Rubin, 1983). This reweights the likelihood of users to be exposed to hate speech (Liang et al., 2016; Schnabel et al., 2016) and we obtain debiased propensity scores. We employ the debiased propensity scores to model the debiased PLV of users to hate speech by learning latent representations in the form of debiased embeddings. 3 We then model the effects of user attributes on the probability to reshare hate speech while controlling for the PLV. Here, we compute the likelihood to reshare each type of hate speech per user by looking at past reshares. Then, we model the (non-linear) effects of the user attributes (verified status, account age, number of posts, followers, and friends) on the user’s probability to reshare hate speech in the future using an explainable boosting machine (EBM) (Nori et al., 2019). [^1] Codes and data for reproducibility are available via an anonymous repository for review and will be available in a public repository upon publication: https://anonymous.4open.science/r/analyzing_user_characteristics_hatespeech_spreaders-2042/README.md
@@ 2.1. Hate speech on social media
Another stream of literature conducted user-level analyses. Targets of hate speech tend to be more popular or high-profile users (ElSherief et al., 2018), while users that post hate speech are characterized by higher social media activity and tend to act in more connected networks (Hua et al., 2020; Ribeiro et al., 2018). When looking at hate speech in online political discourse, people of color of the Democratic Party, white Republicans, and women in general tend to be the main targets (Solovev and Pröllochs, 2022).
@@ 2.2. Inverse propensity scoring
Click data from observational social media data are subject to selection bias. Specifically, previous research identifies two sources of selection bias: self-selection by the users and recommendations from online platforms (Schnabel et al., 2016). On the one hand, users self-select the content they interact with on social media. On the other hand, online platforms tend to recommend only content to users, that a user is likely interested in. This means that click data are biased because the observed interactions were influenced by self-selection and recommendation bias. Hence, we must carefully account for selection bias in our observational click data.
@@ 2.3. Modeling past latent vulnerability
We later transfer the idea of recommendation algorithms to model the PLV of users to hate speech. In particular, we employ recommendation algorithms to learn the debiased PLV of users to hate speech from IPS-reweighted data. By doing so, we represent the PLV as a debiased embedding.
@@ 3.1. Overview
Our objective is to model the effects of user attributes $A$ on the users’ probability to reshare different types of hate speech $Y$ while accounting for the user’s (latent) vulnerability to hate speech $X$. For this, we use BERTopic to cluster posts into different types of hate speech and employ a state-of-the-art, explainable debiasing framework to model the effects.
Input: We input the interactions of users with hate speech posts as a bipartite graph into our framework. Let $\mathcal{U}=\{u_{1},\ldots,u_{i}\}$ denote users and =$\mathcal{H}=\{h_{1},\ldots,h_{j}\}$ denote the hate speech posts. The interaction between user $u\in\mathcal{U}$ and hate speech post $h\in\mathcal{H}$ is then a binary variable denoted as $S_{uh}=\{0,1\}$, where $S_{uh}=1$ if user $u$ shared hate speech post $h$, otherwise $S_{uh}=0$.
@@
Clustering hate speech: In the first step, we employ BERTopic (Grootendorst, 2022) combined with HDBSCAN to cluster hate speech posts $\mathcal{H}$. Subsequently, we label the clusters using automatic labeling with LLAMA-3 (Meta, 2024). Hence, we generate clusters of types of hate speech $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$.
@@
Modeling past latent vulnerability: In the next step, we model the past latent vulnerability $X$ of users to hate speech. For this, we first apply IPS reweighting to debias our observational data, which reweights the propensity $\pi$ of hate speech posts to mimic a randomized trial, and, thereby, we address the issue of selection bias from above. Then, we model the past latent vulnerability to hate speech $X_{u}$ for each user $u$.
@@
Modeling hate speech resharing: In the final step, we model the effects of user attributes $A$ on users’ probability to reshare hate speech $Y$ while controlling for PLV ${X}$. We write the individual user’s probability to reshare hate speech in general as outcome variable $Y_{u}\in[0,1]$ and the probability to reshare hate speech type $c$ as $Y_{uc}\in[0,1]$. Then, we use an explainable boosting machine (EBM) (Nori et al., 2019) to predict the probability to reshare hate speech. For each user $u$ with $m$ user attributes denoted as vector ${a}_{u}=(a_{1},\ldots,a_{m})$, ${a}_{u}\in A$ and PLV embedding ${X}_{u}$ we model the probability to reshare hate speech $Y_{u}$ ($Y_{uc}$ for different types of hate speech $c$).
@@ 3.2. Clustering hate speech (Step 1)
Given hate speech posts $\mathcal{H}$, we employ BERTopic (Grootendorst, 2022), which leverages BERT embeddings and topic modeling, combined with HDBSCAN for clustering. First, we remove stop words, URLs, and emojis. Then, we generate embeddings for each hate speech post using the pre-trained BERT model all-MiniLM-L6-v2 to capture the semantic content. Next, we reduce the embedding dimensionality using UMAP to enhance performance and interpretability. Lastly, we apply HDBSCAN to the reduced embeddings to identify clusters $\mathcal{C}$. HDBSCAN is highly flexible as it does not require pre-specified numbers of clusters and it is suitable for noisy data, which makes it an ideal choice for social media data. In addition, HDBSCAN can deal well with clusters that vary in shape and size (Campello, Ricardo J. G. B. et al., 2013).
Subsequently, we label the clusters using automatic labeling using LLAMA-3 model Llama-3-70b-chat-hf (Meta, 2024). We generate five names per cluster and manually pick the most fitting one by analyzing representative posts. We hereby obtain concise summaries of the primary themes within each cluster, thus providing clear and contextually relevant labels.
3.3. Modeling past latent vulnerability (Step 2)
@@ 3.3.1. IPS reweighting
Given the users $\mathcal{U}$, the hate speech posts $\mathcal{H}$, and the user-hate speech interactions $S_{uh}$, we apply IPS reweighting to alleviate selection bias and obtain debiased propensity scores $\pi$. The propensity score describes the probability that a user is exposed to hate speech.
Assumption: As in (Joachims et al., 2017), we assume that users share posts when (a) they have been exposed to the post, denoted as binary variable $E_{uh}\in\{0,1\}$ ($E_{uh}=1$ if user $u$ was exposed to hate speech post $h$, $E_{uh}=0$, otherwise), and (b) when they are interested in the post, denoted as binary variable $I_{uh}\in\{0,1\}$ ($I_{uh}=1$ if the user is interested in the post, $I_{uh}=0$, otherwise). We further assume that resharing of hate speech is missing-not-at-random (MNAR) (Little and Rubin, 2020). Hence, the probability of user $u\in\mathcal{U}$ spreading a hate speech post $h\in\mathcal{H}$ is a product of the probability of exposure and the probability of interestingness of a post (Saito, 2020). Formally, we have
@@
Why the propensity score is a biased estimate of $\theta_{uh}$: Simply computing $\pi$ from our data gives a biased estimate for $\theta_{uh}$. The reason is that, when working with observational click data, we face two major challenges: First, we only record positive interactions between users and hate speech posts (i.e., hate speech posts that are reshared by the user). Negative interactions (i.e., hate speech posts that were not reshared by the users) are not recorded. Therefore, we do not know whether (a) the user was not interested in a given post or whether (b) the user was not exposed to the post. Second, the observed interactions are likely subject to selection bias. Users are more likely to reshare hate speech posts that are more prevalent, and online platforms are more likely to recommend popular posts to users. Because of the above, the propensity scores $\pi$ is biased when computed from observational social media data, i.e.,
@@
    where we assume that the probability that a user $u$ is exposed to a hate speech post $h$ is higher if the post has gone more viral (as measured by the number of reshares in the bipartite graph). The IPS for rare events could be very high. Hence, we apply a smoothing operator $\mu=0.5$ for more stable propensity scores (see Appendix D for a robustness check with varying $\mu$).
@@
As a result, we obtain debiased propensity scores. Later, we perform an ablation study where we compare the three variants (1)–(3) of our recommendation algorithm to the biased algorithm and find that the debiased recommendation algorithms perform better.
@@ 3.3.2. Modeling past latent vulnerability
Recommender algorithm: In our framework, we employ a Bayesian personalized ranking matrix factorization (BPRMF) (Rendle et al., 2009). We optimize the parameters in the BPRMF using a pairwise Bayesian personalized ranking (BPR) loss (Rendle et al., 2009). By using BPRMF, we address the challenges of missing negative interactions, since it assumes that the observed interactions better explain users’ preferences than unobserved ones and should hence be assigned a higher predictive score. Since we employ three different ways to debias propensity scores, we also obtain three different variants of our debiased recommendation model: BPRMF-V (using $\pi_{\mathrm{virality}}$); BPRMF-F (using $\pi_{\mathrm{follower}}$); and BPRMF-NN (using $\pi_{\mathrm{neural}}$). In Appendix D we perform a robustness check where we compare the performance of our unbiased recommendation models and the biased recommender. We find that the recommendation algorithms with debiased propensity scores are better at predicting the users vulnerability to hate speech. For each user $u$, we model PLV to hate speech as a 64-dimensional embedding ${X}_{u}$. For comparison, the base model BPRMF is biased since it uses $\pi_{\mathrm{biased}}$ as the propensity score.
@@
3.4. Modeling hate speech resharing (Step 3)
3.4.1. Likelihood to reshare hate speech
(12) $Y_{uc}=\frac{n^{u}_{c}}{n^{u}_{\mathrm{hate}}+n^{u}_{\mathrm{normal}}},$